OpenStudy (anonymous):

The sum of two numbers is 57, and the difference is 11. What are these numbers?

OpenStudy (anonymous):

34 and 23

OpenStudy (anonymous):

x+y=57 x-y=11 2x=68 x=34 34-y=11 y=34-11 y=23